OpenStudy (anonymous):

I'm having problems with y'+y=t/(y^2). Whats the procedure?

OpenStudy (athe):

Eq form: \[y^2\cdot y'+y^3=t\] Try use this substitution: \[y=u\cdot v, u=u(t), v=v(t) \rightarrow y'=u'\cdot v + v'\cdot u\] and group: \[v^3(u^2\cdot u'+u^3)+v'\cdot u^3\cdot v^2=t\] \[u^2\cdot u'+u^3=0\], and etc.